# FlowVLA: Thinking in Motion with a Visual Chain of Thought

### v2

@@ Abstract
Many Vision-Language-Action (VLA) models are built upon an internal world model trained via direct next-frame prediction ($v_{t}\rightarrow v_{t+1}$). This paradigm, however, presents a fundamental challenge: it conflates the task of predicting physical motion with that of rendering static appearance, forcing a single mechanism to handle both. This inherent coupling often leads to physically implausible forecasts and inefficient policy learning. To address this limitation, we introduce the Visual Chain of Thought (Visual CoT), a framework that disentangles these processes by compelling the model to first reason about motion dynamics before generating the future frame’s visual appearance. We instantiate this principle by proposing FlowVLA, an autoregressive Transformer that explicitly materializes this reasoning process as $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$”, where $f_{t}$ is an intermediate optical flow prediction. By forcing the model to first commit to a motion plan ($f_{t}$), FlowVLA learns disentangled dynamics, resulting in more coherent visual predictions and significantly more efficient policy learning. Experiments on challenging robotics manipulation benchmarks demonstrate that FlowVLA achieves state-of-the-art performance with substantially improved sample efficiency, pointing toward a more principled foundation for world modeling in VLAs. Project page: https://irpn-lab.github.io/FlowVLA/
@@ 1 Introduction
Figure 1: The FlowVLA Two-Stage Training Paradigm. (Top) Stage 1: World Model Pre-training with Visual CoT. The model learns to predict an intermediate motion representation (Flow T) from an initial frame (Frame T), and then uses both to forecast the subsequent frame (Frame T+k). This iterative process yields physically plausible, long-horizon video predictions. (Bottom) Stage 2: Policy Fine-tuning. The pre-trained world model is adapted for control, fine-tuned to generate precise robot actions (Action T) from visual observations. This paradigm leverages the learned dynamics for efficient and accurate policy learning.
@@
Recent advances in Vision-Language-Action (VLA) models Kim et al. (2024), Zitkovich et al. (2023), Black et al. (2024), Team et al. (2024), particularly those pre-trained as world models like UniVLA Wang et al. (2025) and WorldVLA Cen et al. (2025), have shown remarkable promise for creating generalist robots. The prevailing strategy involves training a large autoregressive transformer to predict the next visual frame given past observations, effectively learning the dynamics of the environment from vast amounts of video data. This learned world model then serves as a powerful foundation for fine-tuning downstream action policies.
@@
Despite their success, these models suffer from a critical, foundational flaw: they conflate the task of physical reasoning with simple pixel prediction. This next-frame prediction paradigm is often a “pixel-copying trap”, where the model learns to replicate static backgrounds without a deep understanding of spatiotemporal dynamics, leading to blurry, inconsistent, and physically implausible long-horizon forecasts Ming et al. (2024). Furthermore, this approach creates a significant domain gap between the passive, observational knowledge learned during pre-training and the active, control-oriented knowledge required for policy learning. This results in inefficient knowledge transfer and requires extensive fine-tuning, as evidenced by slow convergence on downstream tasks Zeng et al. (2024).
@@
We argue that the key to learning physically-grounded dynamics is to force the model to reason explicitly. Drawing inspiration from the success of Chain of Thought (CoT) prompting in Large Language Models Wei et al. (2022), which enhances reasoning by generating intermediate steps, we propose a novel counterpart for world models: a Visual Chain of Thought (Visual CoT). Instead of a single, opaque leap from the current frame $v_{t}$ to the next $v_{t+1}$, we decompose the prediction into a structured reasoning process: first, predict the intermediate physical dynamics—the optical flow $f_{t}$ that describes how every pixel will move. Only then, conditioned on this explicit motion plan, predict the resulting future frame. This $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$ causal chain (see Figure 1 for an overview) transforms the learning objective from mere pattern recognition into a structured physical reasoning task.
@@
We introduce FlowVLA as the first concrete realization of this principle. A key aspect of our design is to integrate motion without introducing dedicated architectural components. We encode optical flow fields as standard RGB-like images, allowing them to be processed by the exact same VQ tokenizer as regular camera observations. This enables a single, unmodified autoregressive transformer to seamlessly learn the interleaved sequence of appearance and motion tokens. This design makes FlowVLA a truly Unified Visual CoT, where the reasoning steps (flow) and states (frames) are expressed in a shared vocabulary and processed by a single, unified model.
@@
  • We identify a fundamental limitation of next-frame prediction for VLA world models and propose Visual Chain of Thought (Visual CoT) as a new principle for learning dynamics.
  • We introduce FlowVLA, a simple yet effective instantiation that unifies appearance and motion reasoning within a single autoregressive Transformer via shared tokenization.
  • We demonstrate through extensive experiments that FlowVLA achieves state-of-the-art performance on challenging manipulation benchmarks, while offering superior sample efficiency, supporting our claim that explicit motion reasoning better bridges the gap between pre-training and policy fine-tuning.
2 FlowVLA
In this section, we introduce FlowVLA, a novel framework designed to instantiate our proposed Visual Chain of Thought (Visual CoT) principle for world model pre-training. We first provide a high-level overview of our two-stage training paradigm. We then detail the core of our contribution: the Visual CoT pre-training stage, including our unified tokenization scheme for appearance and motion. Finally, we describe how the learned world model is finetuned for downstream robotics tasks.
@@ 2.1 Framework Overview
Figure 2: The FlowVLA Framework Architecture. Our model follows a two-stage paradigm. (Left) Stage 1: World Model Pre-training with Visual CoT. Input frames are encoded into appearance tokens (pink). The model then autoregressively predicts an interleaved sequence of motion tokens (blue, representing flow) and future appearance tokens. This structured $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$ prediction forces the model to reason about dynamics before forecasting the future. For conceptual clarity, the Image and Flow Tokenizers are visualized separately; in practice, they are the exact same module applied to both appearance and motion inputs. (Right) Stage 2: Policy Fine-tuning. The pre-trained world model is adapted for control. Conditioned on a text instruction (gray) and the current observation (pink), the model autoregressively predicts action tokens (green) that are decoded into robot actions.
@@
The primary contribution of this paper is concentrated in Stage 1, where we redefine the objective of world model pre-training from simple next-frame prediction to structured physical reasoning.
@@ 2.2 Stage 1: World Model Pre-training via Visual Chain of Thought
A key challenge is to represent two distinct physical signals, appearance (images) and motion (optical flow), within a single model. Our solution is a unified tokenization scheme that preserves architectural simplicity. To process both appearance and motion, we first represent the 2-channel optical flow fields as standard RGB images. Following the technique from VideoJAM Chefer et al. (2025), for each pixel’s flow vector $(u,v)$, we convert it to a color representation based on its polar coordinates. The direction of motion is mapped to the color’s Hue, calculated from the angle $\alpha=\arctan 2(v,u)$. The speed of motion is mapped to the color’s Saturation and Value (Brightness), derived from the vector’s magnitude $m=\sqrt{u^{2}+v^{2}}$. To ensure that subtle movements are not lost while large motions do not saturate the representation, the magnitude is normalized to the range $[0,1]$ using a scaling coefficient $\sigma=0.15$:
@@
where $H$ and $W$ are the height and width of the frame.
Crucially, these resulting flow images, along with the original RGB frames, are processed by the exact same pre-trained VQ-GAN tokenizer Esser et al. (2021). This approach discretizes both modalities into token sequences from a shared vocabulary, offering three key advantages: 1) Parameter Efficiency, as no new motion-specific tokenizer is needed; 2) Architectural Simplicity, maintaining a single, end-to-end autoregressive pipeline; and 3) Unified Representation, enabling the model to learn correlations between appearance and motion in a shared latent space.
@@
$\mathcal{L}_{\text{WM}}=\sum_{t=0}^{T-1}\left(\mathcal{L}_{\text{CE}}(f_{t}|S_{<v_{t+1}})+\lambda\cdot\mathcal{L}_{\text{CE}}(v_{t+1}|S_{<v_{t+1}},f_{t})\right)$ (3)
where $S_{<v_{t+1}}$ denotes all the tokens preceding $v_{t+1}$, and $\lambda$ is a balancing hyperparameter (set to 1.0 in our experiments). This objective explicitly forces the model to perform a “reason $\rightarrow$ predict” process during both training and inference.
2.3 Stage 2: Finetuning for Robotic Control
@@
The policy model is initialized with the weights from the pre-trained FlowVLA. During this stage, the input sequence is composed of interleaved observations and actions: $S_{\text{policy}}=\{L_{\text{instr}},v_{0},a_{0},v_{1},a_{1},\dots\}$, where $a_{t}$ represents the robot’s action tokens.
@@
Actions are discretized into tokens following the FAST Pertsch et al. (2025). Critically, the fine-tuning loss, $\mathcal{L}_{\text{policy}}$, is computed only over the action tokens. This objective directs the model to leverage all its learned visual and dynamical knowledge towards the singular goal of making correct action decisions.
@@ 2. Q2:
    Does explicit motion reasoning lead to superior world modeling capabilities?
@@ 3. Q3:
    Is FlowVLA more sample-efficient during policy finetuning, validating our claim of bridging the pre-training/finetuning gap?
@@ 4. Q4:
    Which components of our design are most critical to its success?
@@ 3.1 Experimental Setup
Benchmarks.
To comprehensively evaluate FlowVLA’s generalization capabilities, we test our method on two distinct and challenging benchmarks: LIBERO and SimplerEnv.
LIBERO Liu et al. (2023) serves as our primary benchmark for evaluating generalization across multiple axes. We follow the standard behavioral cloning setup and report performance on its four main suites, which test generalization to novel spatial layouts, objects, task goals, and long-horizon compositional challenges.
SimplerEnv Li et al. (2024) is used to assess the robustness of our model against significant domain shifts. This benchmark is specifically designed to evaluate policy transfer by introducing diverse variations in lighting, textures, and camera viewpoints, which are more representative of real-world complexity.
@@
Our FlowVLA model is built on the 8.5B parameter Emu3 Wang et al. (2024) and UniVLA Wang et al. (2025) architecture. Our key change is adding optical flow, pre-computed with RAFT Teed and Deng (2020), as an additional input to represent motion. We follow the standard training setup for each benchmark: the model for LIBERO is trained only on the LIBERO dataset, while the model for SimplerEnv is trained only on the Bridge V2 dataset Walke et al. (2023). For LIBERO, we pre-train the world model for 5k steps with a batch size of 16, and then fine-tune the policy for 5k steps with a batch size of 96. For the SimplerEnv benchmark, pre-training runs for 12k steps with a batch size of 32 and policy fine-tuning for 20k steps with a batch size of 128.
3.2 Main Results on Robotics Benchmarks (Q1)
To answer Q1, we evaluate the final performance of FlowVLA after policy finetuning on two distinct and challenging benchmarks: LIBERO and SimplerEnv. Our method establishes a new state-of-the-art on both, demonstrating its effectiveness and robustness.

### v3

# FlowVLA: Visual Chain of Thought-based Motion Reasoning for Vision-Language-Action Models

@@ Abstract
Many Vision-Language-Action (VLA) models are built upon an internal world model trained via next-frame prediction $v_{t}\rightarrow v_{t+1}$”. However, this paradigm attempts to predict the future frame’s appearance directly, without explicitly reasoning about the underlying dynamics. This lack of an explicit motion reasoning step often leads to physically implausible visual forecasts and inefficient policy learning. To address this limitation, we introduce the Visual Chain of Thought, a paradigm that compels the model to first reason about motion dynamics before generating the future frame. We instantiate this paradigm by proposing FlowVLA, an autoregressive Transformer that explicitly materializes this reasoning process as $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$”, where $f_{t}$ is an intermediate optical flow prediction that inherently encodes motion. By forcing the model to first follow the motion plan encoded by $f_{t}$, this process aligns the pre-training objective of dynamics prediction with the downstream task of action generation. We conduct experiments on challenging robot manipulation benchmarks, as well as a real-robot platform. Our FlowVLA not only generates more coherent and physically plausible visual predictions, but also achieves state-of-the-art policy performance with substantially improved sample efficiency, pointing toward a more principled foundation for world modeling in VLAs. Project page: https://irpn-lab.github.io/FlowVLA/
@@ 1 Introduction
Figure 1: Two-Stage Training Paradigm of FlowVLA. (Top) Stage 1: World Model Pre-training with Visual CoT. The model learns to predict an intermediate motion representation (Flow at $T$) from an initial frame (Frame at $T$), and then forecasts the subsequent frame (Frame at T+1{T+1}). This iterative process yields physically plausible, long-horizon video predictions. (Bottom) Stage 2: Policy Fine-tuning. Through fine-tuning, the pre-trained world model is adapted to generate precise robot action chunk (Action at $T$) from visual observations. This paradigm leverages the learned dynamics for efficient and accurate policy learning.
Figure 2: Model Architecture of FlowVLA. Our model instantiates the two-stage training paradigm in Figure 1. (Left) Stage 1: World Model Pre-training with Visual CoT. Input frames are encoded into appearance tokens (pink). The model then autoregressively predicts an interleaved sequence of motion tokens (blue) and future appearance tokens. Our proposed $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$ prediction forces the model to reason about dynamics before forecasting the future. For conceptual clarity, the Image and Flow Tokenizers are visualized separately; in practice, they are the exact same module applied to both appearance and motion inputs. (Right) Stage 2: Policy Fine-tuning. The pre-trained world model is adapted for action prediction. Conditioned on a text instruction (gray) and the current observation (magenta), the model autoregressively predicts action tokens (green) that are decoded into robot action chunk.
@@
Robotics manipulation in diverse and unstructured environments Kim et al. (2024a), Zhou et al. (2025), Xu et al. (2025) has been a long-standing challenge, requiring both precise action prediction and robust understanding of visual and linguistic cues. Recent advances in Vision-Language-Action (VLA) models Kim et al. (2024b), Zitkovich et al. (2023), Black et al. (2024), Team et al. (2024), particularly those pre-trained as world models like UniVLA Wang et al. (2025) and WorldVLA Cen et al. (2025), have shown remarkable promise for creating generalist robots capable of tackling such manipulation tasks. The prevailing strategy involves training a large autoregressive transformer to predict the next visual frame given past observations, effectively learning the dynamics of the environment from vast amounts of video data. This learned world model then serves as a powerful foundation for fine-tuning downstream action policies.
@@
Despite their success, these models suffer from a critical and foundational flaw: they attempt to predict the next frame [^1] Throughout this paper, “frame” denotes a sampled key frame from the video sequence rather than an immediately adjacent frame. in a single, direct step, without explicitly considering the underlying physical motion. This next-frame prediction paradigm is often a “pixel-copying trap” Ming et al. (2024), where the model learns to replicate static backgrounds without a deep understanding of spatiotemporal dynamics, leading to blurry, inconsistent, and physically implausible long-horizon forecasts. Furthermore, this approach creates a significant domain gap between the passive, observational knowledge learned during pre-training and the active, action-oriented knowledge required for policy learning. This results in inefficient knowledge transfer and requires extensive fine-tuning, as evidenced by slow convergence on downstream tasks Zeng et al. (2024).
@@
One of the main reasons for the unsatisfactory performance of the above methods is that they attempt to learn a direct, unreasoned mapping from the current frame to the next, thereby bypassing the crucial step of physical reasoning. Drawing inspiration from the success of Chain of Thought prompting in Large Language Models Wei et al. (2022), which enhances reasoning by generating intermediate steps, we propose a novel paradigm for world models: a Visual Chain of Thought (Visual CoT). Instead of a single, opaque leap from the current frame $v_{t}$ to the next $v_{t+1}$, we decompose the prediction into a structured reasoning process. First, predict the intermediate physical dynamics—the optical flow $f_{t}$ that describes how every pixel will move. Then, conditioned on this explicit motion plan, predict the resulting future frame. This $v_{t}\rightarrow f_{t}\rightarrow v_{t+1}$ causal chain (see Figure 1, top) transforms the learning objective from mere pattern recognition into a structured physical reasoning task. By explicitly modeling dynamics, the world model learns representations that are inherently more aligned with the action-centric knowledge required for policy learning.
To fully leverage the above physically grounded world model for robot action prediction, we propose FlowVLA, a model that realizes the abstract dynamics step using optical flow as the concrete motion representation. Our training paradigm proceeds in two stages as illustrated in Figure 1. Stage 1: World Model Pre-training with Visual CoT focuses on visual reasoning, where the model is pre-trained on large-scale videos to master physically plausible and motion-coherent future frame forecasting. Stage 2: Policy Fine-tuning adapts this pre-trained world model for action prediction: given a text instruction and current observation, the model now predicts discrete robot action chunk instead of future frames. Because Stage 1 has already aligned visual and dynamics representations with physical reality, Stage 2 can fine-tune the policy with substantially improved sample efficiency, directly benefiting from the explicit motion reasoning learned during World Model Pre-training with Visual CoT.
@@
A key aspect of our design is to integrate motion without introducing dedicated architectural components. We encode optical flow fields by a flow map, allowing them to be processed by the exact same Vector Quantization (VQ) tokenizer as regular camera observations. This enables a single, unmodified autoregressive transformer to seamlessly learn the interleaved sequence of appearance and motion tokens. This design makes FlowVLA a truly Unified Visual CoT, where the reasoning steps (flows) and states (frames) are expressed in a shared vocabulary and processed by a single, unified model.
@@
  • We identify a fundamental limitation of next-frame prediction and propose Visual Chain of Thought (Visual CoT) as a new paradigm for learning dynamics for VLA world models, explicitly modeling motion to better bridge the gap between world model pre-training and policy fine-tuning.
  • We introduce FlowVLA, an effective instantiation of this paradigm that unifies appearance and motion reasoning within a single autoregressive Transformer via shared tokenization, avoiding task-specific architectural components and enabling seamless integration of motion reasoning into world model pre-training.
  • We demonstrate through extensive experiments that FlowVLA achieves state-of-the-art performance on challenging manipulation benchmarks and a real-robot platform, while offering superior sample efficiency.
2 Methodology
In this section, we introduce FlowVLA, a novel framework designed to instantiate our proposed Visual Chain of Thought (Visual CoT) paradigm. Firstly, we present our Visual CoT formulation. Secondly, We provide a high-level overview of our two-stage training. Thirdly, we detail the Visual CoT pre-training stage, including our unified tokenization scheme for appearance and motion. Finally, we describe how the learned world model is finetuned for downstream robotics tasks.
2.1 Visual Chain of Thought
The paradigm for pre-training world models is next-frame prediction. This approach aims to learn a probabilistic model, typically parameterized by a large Transformer parameterized by $\theta$, that predicts the next visual observation $v_{t+1}$ given a history of past observations and a guiding language instruction $L$. The learning objective can be expressed as maximizing the log-likelihood of the next frame:
$\max_{\theta}\mathbb{E}_{(v_{t},v_{t+1},L)\sim\mathcal{D}}[\log P_{\theta}(v_{t+1}|v_{t},L)].$ (1)
While conceptually simple and scalable to large and unlabeled video datasets $\mathcal{D}$, this formulation suffers from fundamental limitations that hinder the acquisition of robust physical reasoning.
To overcome the aforementioned challenges, we reframe the world modeling task by introducing a Visual Chain of Thought (Visual CoT). Inspired by the success of CoT in large language models Wei et al. (2022), which improves reasoning by generating intermediate steps, we propose to decompose the visual prediction process. Instead of a single, unreasoned leap, we compel the model to first “think” about the intermediate physical process before generating the final outcome. We instantiate this “thought” as the dense optical flow field $f_{t}$, which describes the per-pixel motion from $v_{t}$ to $v_{t+1}$.
Formally, we reformulate the task from modeling $P(v_{t+1}|v_{t},L)$ in Equation (1) to modeling the joint probability $P(v_{t+1},f_{t}|v_{t},L)$. By applying the chain rule of probability, we factorize this joint probability into a causal sequence:
[…]
This decomposition offers several key advantages. It decouples the learning problem, allowing the model to first focus on the physically-grounded Motion Reasoning task before tackling the more appearance-focused Appearance Generation task. It introduces a powerful inductive bias, explicitly guiding the model to learn a representation of motion, thereby grounding its predictions in physical causality. Crucially, this aligns the pre-training objective with the needs of downstream action prediction. A model that explicitly understands motion $f_{t}$ is inherently better prepared to generate an action chunk $a_{t}$ that cause desired motions.
The above reformulation addresses the limitations of the direct next‑frame prediction paradigm. From a learning perspective, the objective function in Equation (1) is fundamentally ill-posed. It frames world modeling as a high-dimensional regression problem directly in pixel space, which creates an optimization landscape fraught with trivial local minima. This often manifests as the “pixel-copying trap” Ming et al. (2024), where the model discovers that the easiest way to minimize reconstruction error is to simply replicate static background pixels from the input frame. This optimization shortcut is a primary cause of the blurry, inconsistent, and physically implausible forecasts observed in models trained with this paradigm. Furthermore, this direct mapping $v_{t}\rightarrow v_{t+1}$ lacks an explicit causal structure. The model learns a direct correlation between pixel configurations over time, which may not correspond to the true physical causal relationships. This reliance on correlation over causation results in sensitive models that fail to generalize to out-of-distribution scenarios where visual cues change, even if the underlying physics remain the same.
@@ 2.3 Stage 1: World Model Pre-training via Visual Chain of Thought
A cornerstone of FlowVLA’s design is a unified tokenization scheme that represents two physically distinct signals—appearance (images) and motion (optical flow)—within a single, shared vocabulary. This approach preserves architectural simplicity and promotes the learning of deep cross-modal representations. To achieve this, we process each modality as follows.
For the Appearance Representation, standard RGB frames $v_{t}$, which capture the static appearance of the scene, are processed by a pretrained VQ-GAN tokenizer Esser et al. (2021). This tokenizer discretizes each high-resolution image into a compact grid of visual tokens from a learned codebook. This tokenized representation allows the Transformer to process complex visual information in the same sequential format as text.
For the Motion Representation, we encode the abstract motion dynamics using optical flow $f_{t}$. It is a dense pixel-level representation that describes the projected motion of every point in the visual field between two consecutive frames. We choose optical flow for its ability to capture fine-grained interaction dynamics, such as sliding, pushing, and rotating. Another advantage is the availability of robust off-the-shelf models, such as RAFT Teed and Deng (2020), for pre-computation from video data.
We choice optical flow instead of object-centric alternatives like 3D poses or bounding boxes for two main reasons. First, acquiring such object-level representations accurately often requires specialized upstream models trained on large, manually annotated datasets, limiting scalability and introducing potential points of failure. Second, these sparse representations cannot capture non-rigid motion or complex interaction dynamics. In contrast, optical flow provides a dense, general signal that is independent of object detectors. Therefore, it naturally represents continuous motion.
Critically, the image structure of optical flow allows for a shared tokenizer with RGB frames, ensuring tight modality alignment and architectural simplicity. These “flow maps” are then processed by the exact same VQ-GAN tokenizer used for the RGB frames. This design is central to our framework’s efficiency and simplicity, yielding three significant benefits. First, it provides Parameter Efficiency, as no new motion-specific tokenizer or architectural modules are required. Second, it maintains Architectural Simplicity through a single, end-to-end autoregressive pipeline without specialized branches. Finally, it fosters a Unified Representation, encouraging the model to learn deep correlations between appearance (“what is there”) and motion (“how it moves”) within a shared latent space.
To integrate optical flow into our unified framework, we convert the 2-channel flow fields (containing displacements $u$ and $v$) into standard 3-channel RGB images. This conversion, following the technique from VideoJAM Chefer et al. (2025), maps the motion vector at each pixel to a color based on its polar coordinates. Unlike traditional optical flow visualizations that often normalize motion magnitude by the global maximum displacement in a frame, causing subtle motions to be visually suppressed, VideoJAM applies a fixed scaling coefficient and non-linear normalization strategy, which preserves fine-grained motion cues while avoiding saturation in high-speed regions. The direction of motion is mapped to the color’s Hue (from angle $\alpha=\arctan 2(v,u)$), and the speed of motion is mapped to the color’s Saturation and Value (from magnitude $m=\sqrt{u^{2}+v^{2}}$). To handle a wide range of motion speeds without saturation or loss of detail for subtle movements, the magnitude is non-linearly normalized to the range $[0,1]$ using a scaling coefficient $\sigma=0.15$:
@@
where $H$ and $W$ are the frame’s height and width.
@@
[…]
where $S_{<v_{t+1}}$ denotes all the tokens preceding $v_{t+1}$, and $\lambda$ is a balancing hyperparameter (set to 1.0 in our experiments). This objective explicitly forces the model to perform a “reason predict” process during both training and inference.
2.4 Stage 2: Finetuning for Action Prediction
@@
The policy model is initialized with the weights from the pre-trained world model. During this stage, the input sequence is composed of interleaved observations and actions: $S_{\text{policy}}=\{L_{\text{instr}},v_{0},a_{0},v_{1},a_{1},\dots\}$, where $a_{t}$ represents the robot’s action tokens.
@@
Actions are discretized into tokens following the FAST Pertsch et al. (2025). Critically, the fine-tuning loss, $\mathcal{L}_{\text{policy}}$, is computed only over the action tokens. This objective directs the model to leverage all its learned visual and dynamical knowledge towards the singular goal of making correct action prediction.
@@ 2. Q2:
    Does explicit motion reasoning lead to superior world modeling capabilities compared with approaches that learn world dynamics implicitly?
@@ 3. Q3:
    Is FlowVLA more sample-efficient during policy fine-tuning, validating our claim of bridging the pre-training/fine-tuning gap?
@@ 4. Q4:
    What is the effectiveness of each key component in our model architecture?
@@ 3.1 Experimental Setup
To comprehensively assess FlowVLA’s capabilities, we conduct evaluations across a suite of complementary settings. We use two challenging simulation benchmarks, LIBERO and SimplerEnv, to measure generalization and robustness against domain shifts. Additionally, we perform Real-Robot experiments to validate the model’s practical applicability and its ability to transfer skills from simulation to reality.
LIBERO Benchmark. We evaluate FlowVLA on the LIBERO benchmark Liu et al. (2023) to assess its generalization across multiple axes. Following the standard behavioral cloning setup, we report performance on its four main suites, which test for generalization to novel spatial layouts, objects, task goals, and long-horizon compositional challenges.
SimplerEnv Benchmark. We use SimplerEnv Li et al. (2024) to assess the model’s robustness against significant domain shifts. This benchmark is specifically designed to evaluate policy transfer by introducing diverse variations in lighting, textures, and camera viewpoints, which are more representative of real-world complexity.
Real-world Experiments based on AgileX Cobot. As shown in Figure 3(a), we conduct our real-world experiments on a Cobot dual-arm robot manufactured by AgileX Robotics, which adopts the Mobile ALOHA system design Fu et al. (2024). Each arm has 7 degrees of freedom and is equipped with a parallel gripper. The robot carries multiple onboard sensors, including wrist-mounted cameras on both arms and a front-facing camera, for capturing RGB observations used as model input. We design four manipulation tasks to comprehensively evaluate the model’s spatial reasoning and control capabilities (see Figure 3(b)). For each task, we collect 50–200 human-teleoperated demonstrations for fine-tuning, thereby evaluating the model’s data efficiency and its ability to rapidly adapt to the specific embodiment of the Cobot platform.
@@
Our FlowVLA model is built on the 8.5B parameter Emu3 Wang et al. (2024) and UniVLA Wang et al. (2025) architecture. We incorporate optical flow, pre-computed with RAFT Teed and Deng (2020), as an additional modality to represent motion. We follow the standard training setup for the LIBERO and SimplerEnv benchmarks Kim et al. (2024b, 2025). For LIBERO, we pre-train the world model for 5k steps with a batch size of 16, and then fine-tune the policy for 5k steps with a batch size of 96. For the SimplerEnv benchmark, pre-training runs for 12k steps with a batch size of 32 and policy fine-tuning for 20k steps with a batch size of 128.
3.2 Evaluations Results (Q1)
To answer Q1, we evaluate the final performance of FlowVLA after policy finetuning on both benchmarks. Our method establishes a new state-of-the-art on both, demonstrating its effectiveness and robustness.